Question: Find the Complement of the angle measure
ANGLE MEASURE
1.) 37 degrees
2.) 48 degrees
3.) 20 degrees
4.) x degrees
5.) (2x+ 10) degrees
EQUATION
1.) 37+c=90
2.) 48+c=90
3.) 20+c=90
4.) x+c=90
5.) _______
COMPLEMENT
1.) 53 degrees
2.) 42 degrees
3.) 70 degrees
4.) ________
5.) ________

In equation number 1, I just transposed 37 to the right and then I subtract 37 to 90 that's how I got 53 degrees as the complement of the angle measure. But I just couldn't figure out question number 4 and number 5.
There are more Questions but I already answered them I just gave Numbers 1,2,and 3 as examples that you can base on.

Found 2 solutions by MathLover1, richard1234:
Answer by MathLover1(6623)


Question: Find the Complement of the angle measure
ANGLE MEASURE
1.) 37 degrees
2.) 48 degrees
3.) 20 degrees
4.) x degrees
5.) (2x+ 10) degrees
EQUATION
1.) 37+c=90
2.) 48+c=90
3.) 20+c=90
4.) x+c=90
5.) (2x+ 10)=90

COMPLEMENT
1.) 53 degrees
2.) 42 degrees
3.) 70 degrees
4.) __c=90-x ___-------x=0 because it should be added to 90 to get 90
5.) ___2x+ 10=90 _____2x=90-10---2x=80......x=40

In equation number 1, I just transposed 37 to the right and then I subtract 37 to 90 that's how I got 53 degrees as the complement of the angle measure. But I just couldn't figure out question number 4 and number 5.

Answer by richard1234(5390)

Just subtract each angle from 90. The fact that there are variables shouldn't make it any harder.

4: Complement = 90 - x
5: Complement = 90 - (2x + 10) = 80 - 2x